### v1

@@ 4.4 Test 3: Nonlinear resistor-inductor ladder with linear oscillator as signal generator
$y_{r}$ $=h(\pi^{N}(r))$
for our approximate solution $\pi^{N}$ to the invariant mapping. We can compare its steady-state output to the steady-state output of the full-order model. To compare the results, we numerically solve the interconnected system in time for both the full-order model and the reduced-order model, and compare the outputs $y(t)$ and $y_{r}(t)$. In order to simulate these, arbitrary initial conditions must be chosen for $\omega_{0}$ in the signal generating system and for $r_{0}$ in the reduced-order model. Figure 2 shows both outputs for the case $n=1000$, for the mapping $\pi^{N}$ found on the domain $\Omega=[-1,1]^{2}$ with a maximum degree $M=6$, over the time period $t\in[0,10]$. This test was carried out for an initial condition of $\omega_{0}=(0.1,0.2)^{\top}$ and arbitrarily chosen starting points of $r_{0}=(0,1)^{\top}$ and $x_{0}=0$ for the reduced-order model system and the full-order model system, respectively. When constructing the reduced-order model, we use the constant $c=10$. We can see that we are able to recover the full-order model signal using the reduced-order model, and accurately achieve moment matching for this problem.
(a) plot of $y(t)$ vs. $y_{r}(t)$, with $\Omega=[-1,1]^{2}$, $n=1000$.
(b) plot of $|y(t)-y_{r}(t)|$, with $\Omega=[-1,1]^{2}$, $n=1000$.
Figure 2: Plot of $y(t)$ and $y_{r}(t)$ for $t\in[0,50]$ for the RL ladder problem with a linear oscillator signal generator for a domain $\Omega=[-1,1]^{2}$ and a maximum degree $M=6$ for the dimension $n=1000$, with constants $a=2,\kappa=1.1,c=10$ and initial conditions $\omega_{0}=(0.1,0.2)^{\top}$, $r_{0}=(0,1)^{\top}$ and $x_{0}=0$.
We can measure how well the reduced-order model approximates the steady-state response of the system by measuring the RMS error of the response after both the full-order model and reduced-order model signals reach the steady-state. We can then divide this error by the amplitude of the signal to obtain a measure of how well the reduced-order model matches the steady-state response. The RMS error is observed after the simulation reaches steady-state behaviour. Table 7, Table 8, and Table 9 show the relative error for each of the different conditions tested.
We can see that all conditions are able to accurately approximate the steady-state response for $n=2$, $n=100$, and $n=1000$. The performance decreases slightly for the higher dimensional problem, but the difference is not particularly large, and the RMS error is below 0.5 percent for a maximum degree of 6 in all cases, showing that the reduced-order model approximates the full-order model, even when its dimensionality is much less than that of the full-order model.
Domain $M=2$ $M=4$ $M=6$
$\Omega=[-1,1]^{2}$ $3.2250\times 10^{-3}$ $3.5399\times 10^{-4}$ $3.4580\times 10^{-4}$
$\Omega=[-2,2]^{2}$ $1.4806\times 10^{-2}$ $9.6175\times 10^{-4}$ $3.9628\times 10^{-4}$
$\Omega=[-3,3]^{2}$ $3.6235\times 10^{-2}$ $1.7328\times 10^{-3}$ $1.4298\times 10^{-3}$
Table 7: Table of the relative steady-state RMS error of the reduced-order model with $n=2$ for the RL ladder problem with a linear oscillator signal generator with a constants $a=2$ and $\kappa=1.1$.
Domain $M=2$ $M=4$ $M=6$
$\Omega=[-1,1]^{2}$ $3.2222\times 10^{-3}$ $1.5413\times 10^{-3}$ $1.5371\times 10^{-3}$
$\Omega=[-2,2]^{2}$ $1.3348\times 10^{-2}$ $1.9930\times 10^{-3}$ $1.5520\times 10^{-3}$
$\Omega=[-3,3]^{2}$ $3.5621\times 10^{-2}$ $3.2516\times 10^{-3}$ $2.2785\times 10^{-3}$
Table 8: Table of the relative steady-state RMS error of the reduced-order model with $n=100$ for the RL ladder problem with a linear oscillator signal generator with constants $a=2$ and $\kappa=1.1$.
Domain $M=2$ $M=4$ $M=6$
$\Omega=[-1,1]^{2}$ $5.3946\times 10^{-3}$ $4.4702\times 10^{-3}$ $4.4676\times 10^{-3}$
$\Omega=[-2,2]^{2}$ $1.4154\times 10^{-2}$ $4.6679\times 10^{-3}$ $4.4713\times 10^{-3}$
$\Omega=[-3,3]^{2}$ $3.3966\times 10^{-2}$ $5.4216\times 10^{-3}$ $4.7748\times 10^{-3}$
Table 9: Table of the relative steady-state RMS error of the reduced-order model with $n=1000$ for the RL ladder problem with a linear oscillator signal generator with constants $a=2$ and $\kappa=1.1$.
4.5 Test 4: Nonlinear resistor-inductor ladder with Van der Pol oscillator as signal generator
We now consider an example with a nonlinear signal generator. We use the Van der Pol oscillator as the signal generator, modeled as
@@ 4.5 Test 4: Nonlinear resistor-inductor ladder with Van der Pol oscillator as signal generator
The system is the same RL ladder circuit represented in (46). We use the same tensors $\boldsymbol{N}$ and $\boldsymbol{O}$ as in Test 3, and $\boldsymbol{A}$ is now constructed using the new signal generator described by $s(\omega)$ in (54). Using this new $\boldsymbol{A}$ and $Q$, we get the same expressions for $F_{i}$ as shown in (50) and for $JF$ as in (51).
We compute the solution to the invariance equation with a constant value of $\mu=0.25$ and $\kappa=1.1$. Table 10 shows the resulting error $\lVert R(\boldsymbol{c},\omega)\rVert_{(2,W)}$ for the case $n=2$, averaged over the subdomain $W=[-0.7,0.7]^{2}$. Once again, we can see that a higher maximum degree results in a better approximation, but the solution deteriorates as the domain $\Omega$ increases.
Domain $M=2$ $M=4$ $M=6$
$\Omega=[-1,1]^{2}$ $8.0711\times 10^{-3}$ $4.1311\times 10^{-4}$ $2.8741\times 10^{-5}$
$\Omega=[-2,2]^{2}$ $3.9200\times 10^{-2}$ $2.6989\times 10^{-2}$ $7.0562\times 10^{-2}$
$\Omega=[-3,3]^{2}$ $1.0014\times 10^{-1}$ $3.2736\times 10^{-1}$ $1.5181\times 10^{-1}$
Table 10: Table of the average residual error $\lVert R(\boldsymbol{c},\omega)\rVert_{(2,W)}$ over the subdomain $W=[-0.7,0.7]^{2}$ for different values of $\Omega$ and $M$ for the Galerkin approximation to the invariant mapping $\pi$ with $n=2$ for the RL ladder problem with a Van der Pol oscillator signal generator with constants $\mu=0.25$ and $\kappa=1.1$.
Table 11 shows the corresponding results for the higher dimensional $n=100$. For the smaller domains $\Omega=[-1,1]^{2}$ and $\Omega=[-2,2]^{2}$, the results are similar to the previous tests, and the error decreases with a larger maximum degree of the expansion, indicating that the higher maximum degree is better able to approximate the solution to the PDE system. The algorithm failed to converge within 300 iterations for the higher maximum degrees on the domain of $\Omega=[-3,3]^{2}$. Since the signal generator in this example is nonlinear, it is possible that the domain $W^{0}$ where the mapping is valid is smaller than in the case of the linear signal generator; because the nonlinearity contains a monomial of degree 3 in $\mu\omega_{1}^{2}\omega_{2}$, a larger domain will amplify any error near the edges of the domain, and small errors in $\pi^{N}$ will cause large errors in $\lVert R(\boldsymbol{c},\omega)\rVert_{(2,W)}$. The results for $n=1000$ are shown in Table 12. These results are once again very similar to those for $n=100$, and show a good approximation for a domain size up to $\Omega=[-2,2]^{2}$, with the algorithm also failing to converge within 300 iterations on the larger domain.
Domain $M=2$ $M=4$ $M=6$
$\Omega=[-1,1]^{2}$ $6.0786\times 10^{-3}$ $2.9373\times 10^{-4}$ $2.0639\times 10^{-5}$
$\Omega=[-2,2]^{2}$ $5.0803\times 10^{-2}$ $9.7273\times 10^{-3}$ $1.1899\times 10^{-2}$
$\Omega=[-3,3]^{2}$ $4.0017\times 10^{-2}$
Table 11: Table of the average residual error $\lVert R(\boldsymbol{c},\omega)\rVert_{(2,W)}$ over the subdomain $W=[-0.7,0.7]^{2}$ for different values of $\Omega$ and $M$ for the Galerkin approximation to the invariant mapping $\pi$ with $n=100$ for the RL ladder problem with a Van der Pol oscillator signal generator with a constants $\mu=0.25$ and $\kappa=1.1$.
Domain $M=2$ $M=4$ $M=6$
$\Omega=[-1,1]^{2}$ $6.0786\times 10^{-3}$ $2.9373\times 10^{-4}$ $2.0639\times 10^{-5}$
$\Omega=[-2,2]^{2}$ $5.0662\times 10^{-2}$ $4.0561\times 10^{-2}$ $9.0314\times 10^{-3}$
$\Omega=[-3,3]^{2}$ $3.8131\times 10^{-2}$
Table 12: Table of the average residual error $\lVert R(\boldsymbol{c},\omega)\rVert_{(2,W)}$ over the subdomain $W=[-0.7,0.7]^{2}$ for different values of $\Omega$ and $M$ for the Galerkin approximation to the invariant mapping $\pi$ with $n=1000$ for the RL ladder problem with a Van der Pol oscillator signal generator with a constants $\mu=0.25$ and $\kappa=1.1$.
Table 13 shows the amount of computational time required for the algorithm to converge for $\Omega=[-1,1]^{2}$ and $M=6$ for different dimensions.
$n=2$ $n=100$ $n=1000$
time (s) 10.6 438 6955
Table 13: Table showing the computational time, in seconds, required to compute the approximate solution $\pi^{N}$ for $n=2,100,1000$, with $\Omega=[-1,1]^{2}$ and $M=6$.
We once again want to find an expression for the reduced-order model based on the computed $\pi^{N}$. We use a similar approach as we do in the previous example with the linear oscillator signal generator, and this time we use the mapping
(55) $\bar{g}(r,u)=\begin{bmatrix}0\\ \mu(1-r_{1}^{2})+c\end{bmatrix}u$
for some sufficiently large, positive constant $c$. This mapping achieves exponential stability for the resulting reduced-order model, given as
@@
$y_{r}$ $=h(\pi^{N}(r))$
for our approximate solution $\pi^{N}$ to the invariant mapping.
The steady-state output of the reduced-order model is compared to the steady-state output of the full-order model. Figure 3 shows the output of the full-order model in blue and the output of the reduced-order model in red for the case $n=1000$. Similarly to the case of the linear signal generator, the test is carried out for $\Omega=[-1,1]^{2}$ with $M=6$ with initial conditions $\omega_{0}=(0.1,0.2)^{\top}$, $r_{0}=(0,1)^{\top}$ and $x_{0}=0$, over the time period $t\in[0,35]$. We use a constant of $c=10$. In this case, the model is able to recover the steady-state behaviour of the full-order system for $n=1000$. The figures are very similar for $n=2$ and $n=100$, which are not shown.
(a) plot of $y(t)$ vs. $y_{r}(t)$, with $\Omega=[-1,1]^{2}$, $n=1000$.
(b) plot of $|y(t)-y_{r}(t)|$, with $\Omega=[-1,1]^{2}$, $n=1000$.
Figure 3: Plot of $y(t)$ and $y_{r}(t)$ for $t\in[0,50]$ for the RL ladder problem with a Van der Pol oscillator signal generator for a domain $\Omega=[-1,1]^{2}$ and a maximum degree $M=4$ for the dimension $n=1000$, with constants $\mu=0.25,\kappa=1.1,c=10$ and initial conditions $\omega_{0}=(0.1,0.2)^{\top}$, $r_{0}=(0,1)^{\top}$ and $x_{0}=0$.
Once again, we evaluate the reduced-order model approximation by looking at the relative RMS error of the steady-state response. The RMS error is observed after the simulation reaches steady-state behaviour. Table 14, Table 15 and Table 16 show the relative error for each of the different conditions tested.
In the case of the Van der Pol oscillator, we see that for $n=2$ the performance of the reduced-order model is once again quite similar for most of the conditions, and we get a relatively good approximation of the steady-state response. Higher polynomial expansions result in lower RMS error. In the case of $n=100$ and $n=1000$, the higher dimensional expansions all behave quite well for $\Omega=[-1,1]^{2}$ and $\Omega=[-2,2]^{2}$. The RMS error is low in these cases, showing that the reduced-order model remains able to recover the steady-state behaviour of the full-order system in these higher dimensional cases for the Van der Pol oscillator.
Domain $M=2$ $M=4$ $M=6$
$\Omega=[-1,1]^{2}$ $4.5723\times 10^{-2}$ $7.7093\times 10^{-3}$ $1.6723\times 10^{-3}$
$\Omega=[-2,2]^{2}$ $5.1643\times 10^{-2}$ $1.9637\times 10^{-2}$ $4.9413\times 10^{-2}$
$\Omega=[-3,3]^{2}$ $1.2946\times 10^{-1}$ $2.0774\times 10^{-1}$ $5.7344\times 10^{-2}$
@@
Table 14: Table of the relative steady-state RMS error of the reduced-order model with $n=2$ for the RL ladder problem with a Van der Pol oscillator signal generator with constants $\mu=0.25$ and $\kappa=1.1$.
Domain $M=2$ $M=4$ $M=6$
$\Omega=[-1,1]^{2}$ $4.3206\times 10^{-2}$ $7.1173\times 10^{-3}$ $3.4599\times 10^{-3}$
$\Omega=[-2,2]^{2}$ $3.0439\times 10^{-1}$ $1.5848\times 10^{-2}$ $5.5036\times 10^{-3}$
$\Omega=[-3,3]^{2}$ $3.2542\times 10^{-1}$
Table 15: Table of the relative steady-state RMS error of the reduced-order model with $n=100$ for the RL ladder problem with a Van der Pol oscillator signal generator with constants $\mu=0.25$ and $\kappa=1.1$.
Domain $M=2$ $M=4$ $M=6$
$\Omega=[-1,1]^{2}$ $4.3949\times 10^{-2}$ $8.1825\times 10^{-3}$ $5.7352\times 10^{-3}$
$\Omega=[-2,2]^{2}$ $3.0441\times 10^{-1}$ $1.6432\times 10^{-2}$ $6.9633\times 10^{-3}$
$\Omega=[-3,3]^{2}$ $3.2592\times 10^{-1}$
Table 16: Table of the relative steady-state RMS error of the reduced-order model with $n=1000$ for the RL ladder problem with a Van der Pol oscillator signal generator with constants $\mu=0.25$ and $\kappa=1.1$.
5 Concluding remarks
This paper presents a technique for approximating the solution of the invariance PDE system that arises from reduced-order modeling via moment matching for nonlinear systems. These findings for approximating the moment of a system can be applicable to model order reduction problems, as well as immersion and invariance techniques. The solution to the system of PDEs is approximated using a spectral expansion through a chosen set of basis functions. The algorithm to find an approximate solution to this problem can be greatly sped up by constructing matrices and tensors taking advantage of the problem’s structure. Numerical tests are able to find approximate solutions to the invariance PDE system and recover the behaviour of the steady-state responses of the full-order models for nonlinear systems of dimension up to $n=1000$. In cases where an analytical solution exists, the method is able to recover the exact solution in the polynomial basis. In more complex cases, the algorithm can build a reduced-order model that achieves moment matching and produces a good approximation to the steady-state behaviour of the full-order model. ’
A potential future direction to take this research involves using signal generators of higher dimensionality. Currently, we have considered signal generators of dimension $d=2$, which means that our method only involves the computation of two-dimensional integrals in the evaluation of the function $F$. Using higher-dimensional signal generators would allow us to create reduced-order models for problems with more complex input signals, capturing richer dynamical behaviour. In particular, the ability to handle higher-dimensional signal generators could be especially helpful in analyzing multi-input, multi-output systems. Increasing the dimensionality of the signal generator also introduces formidable computational challenges, as leading to large-scale systems of high-dimensional PDEs. In some cases, it may be possible to leverage the problem’s structure in order to compute these, for example if the system’s dynamics are separable.

### v2

@@ 4.2 Test 2: Nonlinear resistor-inductor ladder with linear oscillator as signal generator
$y_{r}$ $=h(\pi^{N}(r)),$
for the approximate solution $\pi^{N}$ to the invariant mapping. Figure 2(a) shows the outputs $y(t)$ and $y_{r}(t)$ for $n=1000$, $M=6$, $\Omega=[-0.6,0.6]^{2}$, $c=10$, and initial conditions $\omega_{0}=(0.2,0.2)^{\top}$, $r_{0}=(0,1)^{\top}$, $x_{0}=0$, over $t\in[0,30]$. The relative RMS error $\epsilon_{rel}$, defined in (48) and reported in Table 2, is on the order of $10^{-3}$ across all values of $M$, confirming that the reduced-order model accurately reproduces the full-order steady-state output despite the large dimensionality reduction.
4.3 Test 3: Nonlinear resistor-inductor ladder with Van der Pol oscillator as signal generator
We now consider the same RL ladder system (51) driven by the Van der Pol oscillator as signal generator,
@@ 4.3 Test 3: Nonlinear resistor-inductor ladder with Van der Pol oscillator as signal generator
The tensors $\boldsymbol{N}$ and $\boldsymbol{O}$ from Test 2 are reused; only $\boldsymbol{A}$ is reconstructed using $s(\omega)$ in (59), and the expressions (55)–(56) for $F_{i}$ and $JF$ remain unchanged.
The invariance equation is solved with $\mu=1$, $\kappa=1.1$, over $\Omega=W=[-0.6,0.6]^{2}$ for $n=1000$. The residual errors in Table 2 confirm that higher polynomial degree yields better approximations, and that the method remains tractable under nonlinear excitation.
For the reduced-order model, we use the mapping
(60) $\bar{g}(r,u)=\begin{bmatrix}0\\ \mu(1-r_{1}^{2})+c\end{bmatrix}u,$
yielding the reduced-order model
@@
$y_{r}$ $=h(\pi^{N}(r)).$
Local exponential stability follows from the linearization of
$\frac{\partial}{\partial r}\left(s(r)-\begin{bmatrix}0\\ \mu(1-r_{1}^{2})+c\end{bmatrix}\ell(r)\right)$
about the origin, which gives the matrix $\begin{bmatrix}0&1\\ -1&-c\end{bmatrix}$, whose eigenvalues have negative real part for all $c>0$.
Figure 2(b) shows the model outputs for $n=1000$, $M=6$, $c=10$, and initial conditions $\omega_{0}=(0.2,0.2)^{\top}$, $r_{0}=(0,1)^{\top}$, $x_{0}=0$, over $t\in[0,30]$. The relative RMS errors $\epsilon_{rel}$ in Table 2 decrease monotonically with $M$, confirming that the polynomial basis has sufficient representational capacity to capture the nonlinear steady-state behaviour induced by the Van der Pol excitation.
$M=2$ $M=4$ $M=6$ $M=8$
$\|R(\mathbf{c},\omega)\|_{(2,W)}$ Test 2 $3.4\times 10^{-4}$ $1.9\times 10^{-6}$ $2.0\times 10^{-8}$ $6.5\times 10^{-10}$
Test 3 $7.4\times 10^{-3}$ $2.0\times 10^{-4}$ $1.2\times 10^{-5}$ $8.1\times 10^{-7}$
$\epsilon_{\mathrm{rel}}$ Test 2 $4.6\times 10^{-3}$ $4.5\times 10^{-3}$ $4.5\times 10^{-3}$ $4.5\times 10^{-3}$
Test 3 $1.7\times 10^{-1}$ $7.0\times 10^{-2}$ $4.9\times 10^{-2}$ $4.1\times 10^{-2}$
@@
Table 2: Galerkin approximation results for the RL ladder system ($n=1000$): residual error $\|R(\mathbf{c},\omega)\|_{(2,W)}$ of the invariant mapping $\pi$ and relative steady-state RMS error $\epsilon_{\mathrm{rel}}$ of the reduced-order model, for Test 2 (linear oscillator) and Test 3 (Van der Pol signal generator).
(a) Test 2: Linear oscillator signal generator.
(b) Test 3: Van der Pol signal generator.
Figure 2: Time evolution of the excitation signal $\omega_{1}(t)$, the full-order output $y(t)$, the reduced-order output $y_{r}(t)$, and the error $|y(t)-y_{r}(t)|$ for the RL ladder system ($n=1000$) over $t\in[0,30]$.
5 Discussion of Results
The numerical results of the previous section demonstrate that the proposed method achieves moment matching and accurately recovers the steady-state output of the full-order model. The residual analysis confirms convergence of the polynomial approximation to the weak solution of the invariance equation with increasing degree. This section provides additional numerical evidence and examines the behaviour of the method under parameter variations, with all experiments based on the RL ladder driven by the Van der Pol oscillator from Section 4.3, with the same constants and initial conditions.
5.1 Effect of domain and nonlinearities
The mapping $\pi$ is only theoretically guaranteed to exist in a neighbourhood $W^{0}$ of the origin, and the polynomial approximation $\pi^{N}$ is constructed over a chosen domain $\Omega$. The size of $\Omega$ therefore directly affects both the approximation quality of $\pi^{N}$ and whether the mapping satisfying (20) is well-defined over that domain. Larger domains amplify higher-order monomials near the boundary and can magnify small coefficient errors.
Figure 3 illustrates this for $\mu=0.25$, $\mu=1$, and $\mu=2$, corresponding to increasingly nonlinear regimes of the Van der Pol signal generator, with residuals measured on $W=[-0.6,0.6]^{2}$. The results confirm that the choice of $\Omega$ has a significant effect on approximation quality, and that increased signal generator nonlinearity heightens sensitivity to the domain. This is attributable to the cubic term $\mu\omega_{1}^{2}\omega_{2}$ in the nonlinearity, which grows rapidly near the boundary when $\Omega$ or $\mu$ are large, thereby amplifying small coefficient errors in the approximate mapping $\pi^{N}$. Despite this, the method maintains stable convergence and acceptable accuracy across the tested parameter range for moderate domain sizes.
Figure 3: PDE residual error as a function of domain size $\Omega$ for different values of the nonlinearity constant $\mu$ in the Van der Pol oscillator signal generator for Test 3. The tests were carried out with a maximum degree $M=6$ for the problem with $n=1000$ dimensions and constants $\kappa=1.1$, $c=10$, and tested on $W=[-0.6,0.6]^{2}$.
5.2 Comparison to other ROM techniques
To provide a benchmark, the proposed method is compared against Proper Orthogonal Decomposition (POD) [42] and Balanced Truncation (BT) [38], two established linear model reduction techniques that do not exploit the structured interconnection or nonlinear excitation present in the considered systems. The implementations are described as follows.
For POD, snapshots of the full-order state $x(t)$ were collected by simulating a single trajectory with an initial condition varying from the one tested, over the time interval $t\in[0,50]$ with initial conditions $\omega_{0}=(0.1,0.3)^{\top}$ and $x_{0}=0$. The snapshot matrix was assembled from 500 uniformly spaced time samples and the POD basis was extracted via a truncated singular value decomposition retaining $r=2$ modes. The reduced dynamics were obtained by Galerkin projection of the full nonlinear system onto the POD subspace.
For BT, the method was applied to the system linearized about the origin. The controllability and observability Gramians were computed analytically by solving the associated Lyapunov equations, and the system was truncated to retain $r=2$ states. The input and output matrices $B$ and $C$ were taken from the linearization of $f$ and $h$, respectively.
For Test 3 with $M=8$ and reduced dimension $r=2$, the proposed method achieves $\epsilon_{rel}=4.1\times 10^{-2}$, compared to $5.7\times 10^{-2}$ for POD and $1.2\times 10^{-1}$ for BT. Figure 4 further illustrates the advantage of moment matching through phase portrait comparisons: the state trajectory reconstructed via $\pi^{N}$ more closely follows that of the full-order model than either POD or BT at the same reduced dimension.
Figure 4: Phase portraits of the full-order model $(x_{1},x_{2})$ and approximate mapping $(\pi_{1}(\omega),\pi_{2}(\omega))$ for Test 3, contrasted with the POD ROM (left) and the BT ROM (right).
5.3 Computational complexity
Table 3 reports the number of basis functions $N$, Newton iterations, and total computation time for Test 3. The Jacobian $JF$ is stored sparsely, so the time complexity scales linearly in $n$ and quadratically in $N$. The rapid convergence of both the residual and RMS error with $M$ means that $M=4$ or $M=6$ already provides a good approximation at manageable cost.
$M=2$ $M=4$ $M=6$ $M=8$
N 55 $14$ $27$ $44$
Iterations 3 3 3 3
Time (s) $1.92$ $80$ $1065$ $8423$
Table 3: Number of basis functions $N$, Newton iterations, and total computational time for the Galerkin approximation of $\pi$ in Test 3.
Regarding online cost, for $t=50$, $n=1000$, and $r=2$, the FOM requires $0.4243$ seconds to simulate while the ROM requires $0.0120$ seconds, a speedup of $35.3\times$. This arises from integrating an $r$-dimensional system rather than the full $n$-dimensional one, with $r\ll n$, with the evaluation of $\pi^{N}$ at the final time adding a cost that scales with $N$. For the offline computation, the tensors $\mathcal{N}$ and $\mathcal{O}$ have sizes $O(N^{3})$ and $O(N^{4})$ respectively; explicit storage was feasible for the problems considered here, though symmetry-based compression or on-demand computation could be employed if storage becomes prohibitive. The Jacobian $JF$ and function evaluation $F$ have sizes $O(N^{2})$ and $O(N)$ respectively, while the coefficients of $\pi^{N}$ require $O(nN)$ storage online.
5.4 Concluding remarks
This paper presents a numerical framework for approximating the solution of the invariance PDE system arising in nonlinear moment matching. A Galerkin expansion through a polynomial basis yields an approximate invariant mapping $\pi^{N}$ that can be used to construct reduced-order models accurately reproducing the steady-state output of systems with state dimension up to $n=1000$. Computational efficiency is achieved by exploiting the problem structure through precomputed matrix and tensor representations.
Nonlinear moment matching is most applicable when the dominant behaviour of a high-dimensional system is governed by a low-dimensional driving subsystem. In this setting, the reduced-order model evolves in a low-dimensional space while the nonlinear complexity of the full-order system is encoded in the mapping $\pi$. Despite the breadth of the theoretical moment matching literature, computational methods for solving the invariance equation in high dimensions have been lacking; this paper addresses that gap directly.
This moment matching method improves upon both POD and BT, and its underlying invariance PDE formulation extends naturally beyond model reduction to output regulation and immersion and invariance design. Future work will address signal generators of higher dimensionality, enabling reduced-order models for richer input regimes while requiring new strategies to manage the associated computational challenges.
6 Code and Data Availability
The code used to generate the numerical results in this paper is publicly available and can be viewed on GitHub at: https://github.com/carlosjdoebeli/Moment-Matching-ROM.